# RAPTOR: Practical Numerical Profiling of Scientific Applications

### v1

@@ Abstract.
The proliferation of low-precision units in modern high-performance architectures increasingly burdens domain scientists. Historically, the choice in HPC was easy: can we get away with 32 bit floating-point operations and lower bandwidth requirements, or is FP64 necessary? Driven by Artificial Intelligence, vendors introduced novel low-precision units for vector and tensor operations, and FP64 capabilities stagnate or are reduced. This is forcing scientists to re-evaluate their codes, but a trivial search-and-replace approach to go from FP64 to FP16 will not suffice.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Arbitrary-precision arithmeticccs: General and reference Empirical studiesccs: Computing methodologies Massively parallel and high-performance simulations
@@
Figure 1. Qualitative visualization of bubble deformation and splitting over time ($t$) for the Bubble test case at Reynolds number $\text{Re}=3500$, performed using incompressible multiphase fluid dynamics solvers in Flash-X. The contours depict the zero level of the level-set function, $\phi=0$, which defines the air-water interface ($\phi>0$ in the air phase and $\phi<0$ in the water phase). Adaptive mesh refinement (AMR) levels $\Omega_{M}$, $\Omega_{M-1}$, and $\Omega_{M-2}$ are shown and labeled in the left panel. AMR dynamically refines the mesh near the interface to accurately capture interface dynamics. We use RAPTOR to truncate FP operations in the advection and diffusion modules of the Navier-Stokes solver following two different strategies. In the first strategy (Trunc. Everywhere), truncation is applied throughout the entire mesh. The second strategy selectively applies truncation in each mesh cell using a cut-off at the $M-l$ AMR level (where $l=1,2,\ldots$). Insets show zoomed-in views of the liquid-gas interface for simulations using 4 bit and 12 bit mantissas, highlighting the effect of reduced precision on the final simulation outcome. Further details in §6.2.
@@
Table 1. Categorization and comparison of RAPTOR to existing approaches and tools. These categories are used: (A) Algorithmic mixed precision, (B) Automatically changing precision, (C) System software-enabled precision changes, (D) Application-granularity, (E) Wrapper and emulator, (F) Precision format, and (G) Observe behavior w/o changing precision. Feature details are as follows: Full app. runs indicates support for unmodified software w/o infeasible runtime increase [^6] Binary for tools which are able to instrument executables.. Dynamic truncation means that truncation of variables can be made algorithmically dependent on the simulation state77footnotemark: 7. Flexible formats stands for arbitrary precision support. Scoped truncation lets the user mark a function/region and the tool truncates the entire call stack below88footnotemark: 8. Granular truncation means that individual operations can be in-/excluded from truncation99footnotemark: 9. Error tracking traces errors through the application code1010footnotemark: 10. Non-$\nabla$ code means that arbitrary (even non-differentiable) code is supported.
@@
languages1111footnotemark: 11
[…]
@@ 1. Introduction
Scientific computing has traditionally relied upon double precision floating-point arithmetic to ensure numerical accuracy and stability across a wide range of problems. However, the growing availability of hardware optimized for lower precision—such as single or even half precision—presents an opportunity to significantly reduce energy consumption and improve computational efficiency. The core dilemma lies in balancing performance gains with the need for reliable scientific results. Some numerical methods are more amenable than others to systematic reduction of precision and can rely on a robust mathematical foundation for maintaining confidence in the outcome. Of these, numerical linear algebra has witnessed the highest level of activity. Several other areas such as many ordinary- and partial-differential equation (ODE and PDE) solvers do not have such mathematically founded support for lowering precision. While many parts of a simulation may tolerate lower precision without compromising accuracy, identifying these regions is nontrivial and highly problem-dependent. Using full precision everywhere guarantees numerical robustness but can be wasteful, whereas an indiscriminate use of reduced precision risks introducing errors that propagate and distort the final outcome as Figure 1 illustrates for a rising bubble simulation.
Here, the incompressible multiphase fluid-dynamics problem involves the deformation and splitting of an air bubble rising in a pool of water. Figure 1 visualizes the evolution of the liquid-gas interface and highlights how different mantissa truncations and numerical precisions (4 bit vs. 12 bit mantissa) affect interface dynamics. A zoomed-in view emphasizes these differences. A full analysis of the experiment is provided in Sections 4 to 6. This example demonstrates that for such workloads the challenge is to develop intelligent strategies and tools that enable selective precision tuning without undermining the fidelity of scientific simulations.
In this paper, we introduce a numerical profiling methodology and accompanying new tool, called RAPTOR, that can numerically profile code execution and enable scientists to infer the impact of reduced precision in an actionable way. We apply RAPTOR to a diverse set of application configurations in Flash-X (Dubey et al., 2022), a multiphysics multidomain application software, where mathematical guidance typically predicts that double precision is needed almost everywhere. We use scientific intuition to selectively reduce precision in code components and observe the impact through the profiles generated by RAPTOR. These experiments effectively demonstrate a methodology where generated insights enable meaningful cost-benefit analysis for running simulation campaigns where mathematics does not provide any guidance.
@@
  • We introduce RAPTOR to easily profile and alter floating-point precision in user-specified code regions for C, C++, and/or Fortran; including MPI support. [Section 3]
@@ 2. Background, Motivation, and Related Work
While mixed-precision arithmetic has seen growing adoption in areas like linear algebra, machine learning, and some PDE solvers, several areas of numerical methods remain relatively underexplored because they involve complex numerical stability and error propagation behavior, making precision tuning less straightforward. Examples include methods such as multigrid (MG) and adaptive mesh refinement (AMR), nonlinear solvers and root-finding algorithms, general-purpose nonlinear solvers, spectral methods used in high-accuracy solutions of differential equations. These methods typically demand high-precision throughout and have not been a focus of mixed-precision research due to the assumption that double precision is required for accuracy preservation.
It is, however, possible to leverage scientific and physical intuition to predict if—and where—it might be feasible to lower the precision. Since there is no robust mathematical framework to validate such assumptions, there is a need for tools that enable interactive and responsive exploration of truncation strategy and provide detailed and immediate insights into the impact of precision changes. We have conducted a review of the existing tools, and define six categories of what and how the approach/tool tackles the numerical analysis and a set of desired features. Table 10 shows these categorizations and feature comparisons, but also highlights that none of the existing tools are feature-complete or easy to apply to an existing, potentially large code base and scientific workloads.
@@ 3. RAPTOR: a Floating-Point Profiling Tool
Our goal for RAPTOR is to numerically profile applications or regions of interest, to determine whether the workload is amenable to execution with lowered precision throughout, under certain conditions only, or not all. We aim to make the iterative process of making changes to an application based on a hypothesis, profiling, and confirming the results as frictionless as possible.
@@ 3.1. Conceptual Design
RAPTOR can scope its transformation to three levels: function, file, and program; and it has two modes of operation: op-mode and mem-mode. The supported configurations of scope and mode are shown in Table 2(a). The mem-mode requires more intervention from the user, and in turn, provides richer information about the computation (cf. Section 3.5 for further details).
@@
RAPTOR consists of two main components: a compiler instrumentation pass and a supporting runtime. The compiler pass inserts calls into the runtime at appropriate places in the user’s program. The runtime executes floating-point operations in the instructed precisions and collects data for analysis. Their place in the compilation pipeline is shown in Figure 2(a).
@@
(a) Supported scopes and features for each mode
Figure 2. Overview of RAPTOR and config. matrix
Figure 3. Usage of RAPTOR to go from original code to either op-mode (right) or mem-mode (left).
Figure 4. RAPTOR’s transformation pass replaces FP operations with runtime calls. We also implement an optimizations which allocates a temporary MPFR variable once to avoid re-allocation for each operation (cf. Figure 5a).
@@ 3.2. Usage
In general, there are flags [^12] The flags differ by compiler/linker and will be explained in the paper’s artifacts. which must be added to load our plugin into the compiler and adjust its optimization pipeline. The simplest way to use RAPTOR is at the file or program scope. This only requires an additional flag for the compilation command, e.g.: --raptor-truncate-all=64_to_5_14;32_to_3_8. The flag instructs RAPTOR to perform a truncation on the currently compiled files for the instructed truncations (in this case 64 bit width floating-point operations to a type with 5 bit exponent and 14 bit mantissa, and analogously for 32 bit width). If specified for all files, we achieve full-program truncation. To use function scope truncation, the code needs to be instrumented as shown in Figure 3. The user can request RAPTOR for a truncated version of a function by using the _raptor_trunc_func_[op,mem] functions, depending on the desired mode. The parameters specify the function and floating-point width to be truncated alongside the target exponent/mantissa lengths [^13] Currently, we require the exponent/mantissa to be compile-time constants.. The returned value is a function with the same type as the input function. Figure 3c shows how the mem-mode truncation can be used. Using mem-mode is more contrived and will be covered in detail in Section 3.5.
@@ 3.3. Instrumentation Pass
The RAPTOR pass is implemented as an instrumentation pass on the LLVM intermediate representation (or LLVM IR) (Lattner and Adve, 2004). LLVM is used by the clang (C/C++) and flang compilers among others. This allows us to handle all of Fortran, C, and C++ under a common representation suitable for compiler transformation and analysis [^14] We are providing examples in C++, however RAPTOR’s concepts apply equivalently to C and Fortran as well (with syntactical differences).. We build on top of Enzyme’s (Moses and Churavy, 2020) infrastructure and make use of its instrumentation and plugin utilities [^15] Enzyme is an auto-differentiation framework for LLVM..
@@
Figure 4a shows an example of the transformation when the user has requested a function scope truncation. It is shown in C++ for illustration purposes, however, the actual code being transformed is in LLVM IR. First, the pass finds all functions that the user requested be truncated and the truncation configuration (in this case, the user requested the function foo to be truncated). Then, we find all transitively called functions (in this case, foo and bar), and for each of them, we replace all operations on floating-point numbers with calls into our runtime. Here, by utilizing features available in LLVM, we can recognize floating-point arithmetic and functions in math libraries, such as the standard C and C++ math library. All affected functions are cloned transformations in order to preserve the behavior of unrelated code that uses them. When the requested scope is file- or program-level, our pass applies the same transformation to the floating-point operations to all functions, without the special handling required for function-scope truncation. This pass is implemented as an LLVM pass plugin, allowing it to be used in compilers such as clang or flang via their plugin interface in supported linkers (lld and gold), and as a standalone pass.
@@
The stage of the optimization pipeline in which the pass runs dictates how closely the truncated operations match the optimized version of the floating-point operation. Compilers have various optimizations that affect floating-point computation, e.g., vectorization or fusion. However, they can only handle operations they are aware of. For example, in Figure 3a, the code contains floating-point arithmetic and a sqrt call to the libc math library, both of which are known to the optimizer, so it can treat them appropriately. However, once the code has been transformed (Figure 4a) into calls into our runtime, the optimizer no longer understands the semantics of the operation and no further floating-point operation optimization is possible.
Ideally, our pass would run after generic floating-point optimizations are done to preserve as many compiler optimizations as possible. However, due to current limitations (see Section 7.3), we are not able to handle vectorized instructions, which means we opt to run our pass before the vectorization passes.
@@
If a call to a function with no definition is encountered while attempting to truncate, we are not able to instrument any operations inside it. For this reason, we configure the compiler(s) to use link-time optimizations (LTO), which merges all LLVM modules from different files into one file. We run our pass after the merge, which allows us to analyze the entire call graph [^16] Calls to pre-compiled external libraries will be ignored and we emit a warning., see Figure 2(a).

### v2

@@ Abstract.
The proliferation of low-precision units in modern high-performance architectures increasingly burdens domain scientists. Historically, the choice in HPC was easy: can we get away with 32 bit floating-point operations and lower bandwidth requirements, or is FP64 necessary? Driven by Artificial Intelligence, vendors introduce novel low-precision units for vector and tensor operations, and FP64 capabilities stagnate or are reduced. This forces scientists to re-evaluate their codes, but a trivial search-and-replace approach to go from FP64 to FP16 will not suffice.
@@
Mixed precision, low precision, numerical profiling, error tracking, simulation accuracy, multiphysics, LLVM, MPFR
journalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759810isbn: 979-8-4007-1466-5/2025/11ccs: Mathematics of computing Arbitrary-precision arithmeticccs: General and reference Empirical studiesccs: Computing methodologies Massively parallel and high-performance simulations
@@
Figure 1. Qualitative visualization of bubble deformation and splitting over time ($t$) for the Bubble test case at Reynolds number $\text{Re}=3500$, performed using incompressible multiphase fluid dynamics solvers in Flash-X. The contours depict the zero level of the level-set function, $\phi=0$, which defines the air-water interface ($\phi>0$ in the air phase and $\phi<0$ in the water phase). Adaptive mesh refinement (AMR) levels $\Omega_{M}$, $\Omega_{M-1}$, and $\Omega_{M-2}$ are shown and labeled in the left panel. AMR dynamically refines the mesh near the interface to accurately capture interface dynamics. We use RAPTOR to truncate FP operations in the advection and diffusion modules of the Navier-Stokes solver following two different strategies. In the first strategy (Trunc. Everywhere), truncation is applied throughout the entire mesh. The second strategy selectively applies truncation in each mesh cell using a cut-off at the MlM-l AMR level (where $l=1,2,\ldots$). Insets show zoomed-in views of the liquid-gas interface for simulations using 4 bit and 12 bit mantissas, highlighting the effect of reduced precision on the final simulation outcome. Further details in §6.2.:
@@
Table 1. Categorization and comparison of RAPTOR to existing approaches and tools. These categories are used: (A) Algorithmic mixed precision, (B) Automatically changing precision, (C) System software-enabled precision changes, (D) Application-granularity, (E) Wrapper and emulator, (F) Precision format, and (G) Observe behavior w/o changing precision. Feature details are as follows: Full app. truncation indicates support for unmodified software w/o infeasible runtime increase. [^2] E.g., Flash-X can not realistically be analyzed with ADAPT without applying kernel extraction or similar techniques to isolate individual regions of interest in the workload. Dynamic truncation means that truncation of variables can be made algorithmically dependent on the simulation state. [^3] E.g., NEAT has functionality to change behavior based on the calling context. Flexible formats stands for arbitrary precision support. Scoped truncation lets the user mark a function/region and the tool truncates the entire call stack below. [^4] E.g., tools making use of LLVM IR have this functionality. Granular truncation means that individual operations can be in-/excluded from truncation. [^5] E.g., the approach by Gu et al. considered individual flops for tuning. Error tracking traces errors through the application code. [^6] E.g., GPUMixer demonstrated this by using shadow variables. Non-$\nabla$ Code means that arbitrary (even non-differentiable) code is supported.
@@
languages [^7] Binary for tools which are able to instrument executables.
ADAPT B , , Σ , Σ Σ , C, C++, Fortran (Menon et al., 2018)
CADNA C Σ , , N/A Σ Σ Σ Ada, C, Fortran (Jézéquel et al., 2021)
FPSpy G Σ , , N/A N/A , Σ Binary (Dinda et al., 2020)
FPVM D Σ , Σ , , Σ Σ Binary (Dinda et al., 2022)
GPU-FPX G Σ , N/A N/A , Σ Σ GPU Binary (Li et al., 2023)
GPUMixer B , , , , , Σ , CUDA (Laguna et al., 2019)
Jost et al. C, E, F Σ , Σ N/A Σ , Σ C (Jost et al., 2021)
Gu et al. E , , , , Σ , Σ C, C++ (Gu et al., 2020)
NEAT E Σ Σ Σ , Σ , Σ Binary (Barati et al., 2021)
Precimonious A , N/A , Σ , , Σ C (Rubio-González et al., 2013)
Puppeteer D Σ Σ N/A , Σ , , C, C++ (Parasyris et al., 2022)
RAPTOR B, C, E Σ Σ Σ Σ Σ Σ Σ C, C++, Fortran
@@ 1. Introduction
Scientific computing has traditionally relied on double precision floating-point arithmetic to ensure numerical accuracy and stability across a wide range of problems. However, the growing availability of hardware optimized for lower precision—such as single or even half precision—presents an opportunity to significantly reduce energy consumption and improve computational efficiency. The core dilemma lies in balancing performance gains with the need for reliable scientific results. Some numerical methods are more amenable than others to systematic reduction of precision and can rely on a robust mathematical foundation for maintaining confidence in the outcome. Of these, numerical linear algebra has witnessed the highest level of activity. Several other areas such as many ordinary- and partial-differential equation (ODE and PDE) solvers do not have such mathematically founded support for lowering precision. While many parts of a simulation may tolerate lower precision without compromising accuracy, identifying these regions is nontrivial and highly problem-dependent. Using full precision everywhere guarantees numerical robustness but can be wasteful, whereas an indiscriminate use of reduced precision risks introducing errors that propagate and distort the final outcome. Figure 1 illustrates this for a rising bubble simulation.
Here, the incompressible multiphase fluid-dynamics problem involves the deformation and splitting of an air bubble rising in a pool of water. Figure 1 visualizes the evolution of the liquid-gas interface and highlights how different mantissa truncations and numerical precisions (4 bit vs. 12 bit mantissa) affect interface dynamics. Zoomed-in views emphasize these differences. A full analysis of the experiment is provided in Sections 4 to 6. This example demonstrates that for such workloads the challenge is to develop intelligent strategies and tools that enable selective precision tuning without undermining the fidelity of scientific simulations.
In this paper, we introduce a numerical profiling methodology and accompanying new tool, called RAPTOR, that can numerically profile code execution and enable scientists to infer the impact of reduced precision in an actionable way. We apply RAPTOR to a diverse set of application configurations in Flash-X (Dubey et al., 2022), a multiphysics multidomain application software, where mathematical guidance typically predicts that double precision is needed almost everywhere. We use scientific intuition to selectively reduce precision in code components and observe the impact through the profiles generated with the help of RAPTOR. These experiments effectively demonstrate a methodology where generated insights enable meaningful cost-benefit analysis for running simulation campaigns where mathematics does not provide any guidance.
@@
  • We introduce RAPTOR [^1] https://github.com/RIKEN-RCCS/RAPTOR to easily profile and alter floating-point precision in user-specified code regions for C, C++, and Fortran; including GPU and OpenMP support. [Section 3]
@@ 2. Background, Motivation, and Related Work
While mixed-precision arithmetic has seen growing adoption in areas like linear algebra, machine learning, and some PDE solvers, several areas of numerical methods remain relatively underexplored because they involve complex numerical stability and error propagation behavior, making precision tuning less straightforward. Examples include methods such as multigrid (MG) and adaptive mesh refinement (AMR), nonlinear solvers and root-finding algorithms, general-purpose nonlinear solvers, spectral methods used in high-accuracy solutions of differential equations. These methods typically demand high precision throughout and have not been a focus of mixed-precision research due to the assumption that double precision is required for accuracy preservation.
It is, however, possible to leverage scientific and physical intuition to predict if—and where—it might be feasible to lower the precision. Since there is no robust mathematical framework to validate such assumptions, there is a need for tools that enable interactive and responsive exploration of truncation strategies and provide detailed and immediate insights into the impact of precision changes. We have conducted a review of the existing tools, and define six categories of what and how the approach/tool tackles the numerical analysis and a set of desired features. Table 6 shows these categorizations and feature comparisons, but also highlights that none of the existing tools are feature-complete or easy to apply to an existing, potentially large code base and scientific workloads.
@@ 3. RAPTOR: a Floating-Point Profiling Tool
Our goal for RAPTOR is to numerically profile applications or regions of interest to determine whether the workload is amenable to execution with lowered precision throughout, under certain conditions only, or not all. We aim to make the iterative process of making changes to an application based on a hypothesis, profiling, and confirming the results as frictionless as possible.
@@ 3.1. Conceptual Design
RAPTORcan scope its transformation to three levels: function, file, and program; and it has two modes of operation: op-mode and mem-mode. The supported configurations of scope and mode are shown in Figure 2(b). The mem-mode requires more intervention from the user, and in turn, provides richer information about the computation (cf. Section 3.5 for further details).
@@
RAPTORconsists of two main components: a compiler instrumentation pass and a supporting runtime. The compiler pass inserts calls to the runtime at appropriate places in the user’s program. The runtime executes floating-point operations in the instructed precisions and collects data for analysis. The placement of the components in the compilation pipeline is shown in Figure 2(a).
@@
(b) Supported scopes and features for each mode
Figure 2. Overview of RAPTOR and config. matrix:
Figure 3. Usage of RAPTOR to go from original code to either op-mode (right) or mem-mode (left).:
{subcaptiongroup}
[…]
Figure 4. RAPTOR’s transformation pass replaces FP operations with runtime calls. We also implement an optimization which allocates a temporary MPFR variable once to avoid re-allocation for each operation (cf. Figure 5). :
{subcaptiongroup}
[…]
Figure 5. Implementation of the RAPTOR runtime.:
@@ 3.2. Usage
In general, there are flags [^8] The flags differ by compiler/linker and are detailed in the paper’s artifacts. which must be added to load our plugin into the compiler and adjust its optimization pipeline. The simplest way to use RAPTOR is at the file or program scope. This only requires an additional flag for the compilation command, e.g.: --raptor-truncate-all=64_to_5_14;32_to_3_8. The flag instructs RAPTOR to perform a truncation on the currently compiled files for the instructed truncations (in this case 64 bit width floating-point operations to a type with 5 bit exponent and 14 bit mantissa, and analogously for 32 bit width). If specified for all files, we achieve full-program truncation. To use function scope truncation, the code needs to be instrumented as shown in Figure 3. The user can request a truncated version of a function using the _raptor_trunc_func_[op,mem] functions, depending on the desired mode. The parameters specify the function and floating-point width to be truncated alongside the target exponent/mantissa lengths. [^9] Currently, we require the exponent/mantissa to be compile-time constants. The return value is a function with the same type as the input function. Using mem-mode truncation, shown in Figure 3c, is more contrived and will be covered in detail in Section 3.5.
@@ 3.3. Instrumentation Pass
The RAPTOR pass is implemented as an instrumentation pass on the LLVM intermediate representation (LLVM IR) (Lattner and Adve, 2004). LLVM is used by the clang (C/C++) and flang compilers among others. This allows us to handle all of Fortran, C, and C++ under a common representation suitable for compiler transformation and analysis. [^10] We are providing examples in C++, however RAPTOR’s concepts apply equivalently to C and Fortran as well (with syntactical differences). We build on top of Enzyme’s (Moses and Churavy, 2020) infrastructure [^11] Enzyme is an auto-differentiation framework for LLVM. and make use of its instrumentation and plugin utilities.
@@
Figure 4 shows an example of the transformation when the user has requested a function scope truncation. It is shown in C++ for illustrative purposes, but the actual code being transformed is in LLVM IR. First, the pass finds all functions that the user requested be truncated and the truncation configuration (in this case, the user requested the function foo to be truncated). Then, the pass finds all transitively called functions (in this case, foo and bar), and for each of them, replaces all operations on floating-point numbers with calls into the RAPTOR runtime. By utilizing features available in LLVM, we can recognize floating-point arithmetic and functions in math libraries, such as the standard C and C++ math library. All affected functions are cloned transformations in order to preserve the behavior of unrelated code that uses them. When the requested scope is file or program-level, our pass applies the same transformation to the floating-point operations of all functions, without the special handling required for function-scope truncation. This is implemented as an LLVM pass plugin, which can be used in compilers such as clang or flang via their plugin interface in supported linkers (lld and gold), and as a standalone pass.
@@
The stage of the optimization pipeline in which the pass runs dictates how closely the truncated operations match the optimized version of the floating-point operation. Compilers have various optimizations that affect floating-point computation, e.g., vectorization or fusion. However, they can only handle operations they are aware of. For example, in Figure 3a, the code contains floating-point arithmetic and a sqrt call to the libc math library, both of which are known to the optimizer, so it can treat them appropriately. However, once the code has been transformed to call into our runtime (Figure 4), the optimizer no longer understands the semantics of the operation and no further floating-point operation optimization is possible.
Ideally, our pass would run after all generic floating-point optimizations are done to preserve as many compiler optimizations as possible. However, due to current limitations (see Section 7.3), we are not able to handle vectorized instructions, which means we opt to run our pass before the vectorization passes.
@@
If a call to a function with no definition is encountered while attempting to truncate, we are not able to instrument any operations inside it. For this reason, we configure the compiler(s) to use link-time optimizations (LTO), which merges all LLVM modules from different files into one file. We run our pass after the merge, which allows us to analyze the entire call graph, [^12] Calls to pre-compiled external libraries are ignored and RAPTOR emits a warning. see Figure 2(a).